OpenStudy (anonymous):

How do you simplify -5+i/2i

OpenStudy (anonymous):

5i/2i?

OpenStudy (anonymous):

2i^2? (I suck at math so...)

OpenStudy (phi):

yes. and i^2 is special.... it also means -1

OpenStudy (anonymous):

So all in all it's 5i/2i^2?

OpenStudy (phi):

the bottom 2i^2 which can be rewritten as 2* -1= -2 the top is (-5+i)* i which can be simplified (distribute the i)

OpenStudy (anonymous):

-6 right?

OpenStudy (anonymous):

or actually 6

OpenStudy (phi):

i*(-5+i) is done the same way as x(a+b) = ax+bx you should get i*-5 + i*i can you simplify that?

OpenStudy (anonymous):

4?

OpenStudy (phi):

you can't ignore the i, it has to stay ....

OpenStudy (anonymous):

*5

OpenStudy (anonymous):

oh then its 6i? (-1)-1=1 and -5(-1)=5 and 5+1=6i

OpenStudy (anonymous):

So 6i/-2i

OpenStudy (phi):

let's start at the beginning \[ \frac{ (-5+i)}{2i} \] people don't like i in the denominator. so multiply top and bottom by i \[ \frac{ (-5+i)}{2i} \cdot\frac{i}{i}\] what is the bottom? (what is 2*i*i ?)

OpenStudy (anonymous):

2

OpenStudy (phi):

almost. remember i*i is -1 try again

OpenStudy (anonymous):

2*-1=-2. -2*-1=2. and -1*-1=1, 2*1=2.

OpenStudy (phi):

all of that is true. but what is 2*i*i ?

OpenStudy (phi):

the way you do it is bit by bit 2*i*i is 2 * (i*i) replace i*i with -1 2 * -1 replace 2*-1 with -2 so the bottom is -2

OpenStudy (phi):

you now have (after replace 2*i*i with -2) \[ \frac{ (-5+i)i}{-2} \]

OpenStudy (anonymous):

But wouldn't i*i=-1*-1 is i=-1

OpenStudy (phi):

i*i= -1 end of story

OpenStudy (anonymous):

alright and we can solve the numerator normally?

OpenStudy (phi):

i is sqrt(-1) and sqrt(-1) * sqrt(-1) is -1 just remember i*i= -1

OpenStudy (phi):

now for the numberator. first step is "distribute" the i can you do that?

OpenStudy (anonymous):

Not sure how I would.

OpenStudy (phi):

distribute is a rule you should learn here is the rule with numbers (not letters) 2*(3+4) we know this is 2*7 (right?) = 14 notice that we can do this 2*3 +2*4 6+8 14 here is the rule with letters: x*(a+b) x*a + b*x you multiply the "thing" on the outside times each thing on the inside try it with i * (-5 + i)

OpenStudy (anonymous):

So its i*i+i*-5?

OpenStudy (phi):

yes. so now we have \[ \frac{i*i + i* -5}{-2} \] now can you replace i*i?

OpenStudy (anonymous):

-1+5 right?

OpenStudy (phi):

i*i is -1 but you don't drop i from -5*i so far we have \[ \frac{-1 +i*-5}{-2} \]

OpenStudy (phi):

to make it look nicer we don't have to write the * (for multiply) we can just write -5i (which means -5*i which is the same as i*-5 , you can swap order when multiplying) so we now have \[ \frac{-1 -5i}{-2} \] I would multiply top and bottom by -1 can you do that ?

OpenStudy (anonymous):

5i/2?

OpenStudy (phi):

you did the bottom correctly now re-do the top: -1*(-1-5i) remember the "distributive law" (see up above where we used it)

OpenStudy (anonymous):

So 6i? -1*-1+-5*-1?

OpenStudy (phi):

-1*-1+-5*-1? is almost correct except I don't see i in there

OpenStudy (anonymous):

Yea sorry, was called away. But thanks.

OpenStudy (phi):

you are almost done here. -1*(-1-5i) you multiply the "thing" on the outside times each thing on the inside (and remember to keep the i.